# Quantum adiabatic optimization with Rydberg arrays: localization phenomena and encoding strategies

### v1

@@ E.1 Matrix-transfer method and bound state
Figure 17: Analysis of the bound state in a infinite $L$-shaped domain. (a) For $L\to\infty$, the domain of the Helmholtz equation in Eq. (6) acquires an infinite $L$ shape. The two leads are denoted by $1$ and $2$, with widths $\zeta_{1}$ and $\zeta_{2}$, and the inner region with “In”. Three different reference frames are used for the three regions. (b) Ground-state energy, $k^{2}_{\rm GS}$, computed via (black) finite element method (FEM) and (orange) single-mode approximation (SMA) as a function of $r_{\zeta}=\zeta_{2}/\zeta_{1}$. The gap $d$ in Eq. (42) computed using FEM is shown in blue. Here, we set $L=40\zeta_{1}$.
Here, we generalize the matrix-transfer method presented in Ref. [82] to solve the Helmholtz equation with homogeneous DBC in a symmetric infinite $L$-shaped domain, to the asymmetric case (namely, leads of different widths). Following the notations of Fig. 17(a), we divide the Hilbert space into three distinct regions: two leads, denoted by $\Gamma\in\{1,2\}$, of different widths ($\zeta_{1}\equiv\zeta_{x}$ and $\zeta_{2}\equiv\zeta_{y}$); and an inner region, denoted by “In”. We employ three different reference frames: $\{x_{1},y_{1}\}$ in the lead extending along $y$, $\{x_{2},y_{2}\}$ in the lead extending along $x$, and $\{X,Y\}$ in the inner region.
@@
This equation implies ${M_{11}=0}$, from which we determine the value of $k$ for the bound state, $k_{\rm BS}$. We re-scale the system by setting (without loss of generality) $\zeta_{1}=1$ and re-scaling the energies accordingly. We find that $k_{\rm BS}$ satisfy the following equation:
@@ E.2 Energy gap scaling for arbitrary geometries
Here, we employ a finite element method (FEM) to solve the Helmholtz equation with homogeneous DBC, in the finite $L$-shaped domain. For different values of $r_{\zeta}$, we compute how the energy gap ${\delta E=t\left(k^{2}_{\rm th}-k^{2}_{\rm GS}\right)}$ scales with $L$. We choose, without loss of generality, $L=1$ and numerically solve the equation re-scaling the frame of coordinates as $\left\{x^{\prime}=x/L,y^{\prime}=y/L\right\}$ and the energies via ${k^{\prime}=k/L}$. We find the following scaling of the energy gap:
@@ E.3 Conformal mapping and bound state
The bound state solution of the Helmholtz equation can be further investigated via a conformal mapping of the $L$-shaped domain into a unit disk. Here, the transformed equation admits a semi-analytical treatment, which allows us to classify the eigenstates into bound states and propagating states at infinity. Without loss of generality, in the following discussion, we normalize the width of the two leads such that $\zeta_{1}\zeta_{2}=1$.
We parameterize the $L$-shaped domain $\mathcal{L}$ with a complex coordinate $w=\xi+i\eta$, and the unit disk ${\mathcal{D}=\{(x,y)|\sqrt{x^{2}+y^{2}}<1\}}$ with $z=x+iy$. We derive a conformal mapping between $w$ and $z$ in two steps, where we resort to an intermediate coordinate $u$ spanning the upper half of the complex plane. First, the mapping from $u$ to $w$ is achieved via a Schwarz–Christoffel transformation that sends the prevertices $u=-\zeta_{1}^{2},0,\zeta_{1}^{-2},\infty$ on the real axis $u$ to the vertices of the $L$-shaped polygon $w=+i\infty,0,+\infty,\zeta_{1}+i\zeta_{1}^{-1}$, respectively:
@@ E.3.1 Qualitative analysis of the eigenstates
The asymptotic behaviors of the eigenstates at infinity of $\mathcal{L}_{p}$ can be inferred from the analytical solution of Eq. (45) in the vicinity of the transformed points. The points at infinity in the two leads $\Gamma\in\{1,2\}$, namely, $w_{1}=-e^{-ip\pi}\infty$ and $w_{2}=+\infty$, are respectively mapped to $z_{1}=-e^{-i\varphi}$ and $z_{2}=e^{-i\varphi}$ on the unit disk. In the vicinity of these poles, Eq. (45) can be solved by separating the variables after switching to polar coordinates, i.e., $z-z_{\Gamma}=\rho e^{i\theta}$. To the lowest order in $\rho$, it results that $\chi_{p}\simeq\pi^{2}\zeta_{\Gamma}^{4p}/\rho^{2}$ has no angular dependence. By setting $\phi_{p}(\rho,\theta)=R(\rho)T(\theta)$, Eq. (45) returns ${T^{\prime\prime}+m^{2}T=0}$ and
@@
Here, the first inequality is derived from the spatial symmetries of $\phi_{p=0}$ and $\chi_{p}$. Indeed, the invariance of the operator $-\chi_{0}^{-1}\nabla^{2}$ upon the transformation $(x,y)\to(-x,-y)$ leads to $\phi_{p=0}^{2}(x,y)=\phi_{p=0}^{2}(-x,-y)$, and the Jacobians verify ${\chi_{p}(x,y)+\chi_{p}(-x,-y)>2\chi_{p=0}(x,y)}$ for any $p>0$ (using $\mu+1/\mu>2$ for $\mu>0,\mu\neq 1$). The second inequality stems from the variational principle for estimating the ground-state energy (i.e., $E_{p}[\phi]\geq E_{p}[\phi_{p}]$ for every $\phi$). Solving the Helmholtz equation directly in $\mathcal{L}_{0}$ results in $\phi_{p=0}=\frac{1-x^{2}-y^{2}}{\sqrt{[x^{2}+(y-1)^{2}][x^{2}+(y+1)^{2}]}}$ and $k_{p=0}^{2}=\pi^{2}$, and thus the ground state energy in $\mathcal{L}_{p>0}$ is characterized by a strict upper bound: $k_{p>0}^{2}<\pi^{2}$.
@@
E.3.2 Semi-analytical solution
We now present a possible semi-analytical treatment of the ground state for $\mathcal{L}_{p}$ domains. In particular, the homotopy perturbation method (HPM) [90] can be applied to Eq. (45), leading to a recursive approximation of the bound state. Starting from an initial guess $\phi_{p,n=0}$, this method generates a sequence of un-normalized wavefunctions $\phi_{p,n}$ and corresponding energies $k^{2}_{p,n}$, via the following Poisson equation:
@@ Appendix F Analyses beyond the blockade limit
Here, we discuss how the analyses conducted in the blockade limit also hold when considering the Van der Waals interaction $U_{v,u}$ in Eq. (2) beyond the blockade radius. Specifically, we focus on the domain-wall localization phenomenon occurring in the CWE setup (see Secs. IVC and V), but we mention that analogous reasoning applies to the vertex-wire and the crossing gadget.
Due to the tail interaction among non-blockaded atoms, the effective Hamiltonian in Eq. (5) acquires additional site-dependent diagonal terms. Specifically, the diagonal term $\epsilon(\bm{j})$ is given by the sum of the detuning $\Delta_{v}$ on the excited atoms of the domain-wall state $|\bm{j}\rangle$, plus the sum of the interactions $U_{v,u}$ between its pairs of excited atoms (at a distance $d_{v,u}$). In the case of the CWE setup of Fig. 1(d), due to its geometrical structure, these diagonal terms tend to localize the domain walls around the gadget. Indeed, to enforce the blockade constraints, the gadget atoms need to be placed at a shorter distance than those in the legs. This introduces an energetic penalty for states with more excited atoms near the gadget.
In the following, we focus on two possible strategies to mitigate this effect and recover the blockade-limit ground state with high fidelity. The first strategy relies on adjusting the position of the atoms along the vertex-wires, while the second one relies on adjusting the local detuning on each atom. In both strategies, we aim to minimize the spreading of the diagonal terms $\delta\epsilon\equiv\max_{\bm{j}}{\left[\epsilon(\bm{j})\right]}-\min_{\bm{j}}{\left[\epsilon(\bm{j})\right]}$, since the blockade-limit is recovered in the $\delta\epsilon\ll t$ limit.
In the numerical simulations, we consider the parameters of Sec. VA: $\Omega=\Delta/3=7.5$ rad/µs, $C_{6}=5.42\times 10^{6}$ µm6 $\times$ rad/µs, $R_{b}=\left(C_{6}/\sqrt{(2\Omega)^{2}+\Delta^{2}}\right)^{1/6}=7.65$ µm, and the constraint of $d_{v,u}>4$ µm; include the tail interaction between non-blockade atoms; and employ the exact diagonalization method. In the implementation of Sec. VA, we adopted the position-based mitigation strategy.
@@ F.1 Position optimization to mitigate the tail interaction
Here, we optimize the position of the vertex-wire atoms to minimize $\delta\epsilon$, while fixing the position of the five atoms at the intersection of the two vertex-wires. In particular, we set the distance between $G_{c}$ and its neighbors [see inset in Fig. 1(d)] to $0.9R_{b}/\sqrt{2}$ and the other distances (e.g., between $G_{x}$ and $G_{y}$) to $0.9R_{b}$. Then, we search for an optimal position pattern along the vertex-wires, constraining the distance between neighboring atoms to be smaller than $6.880.9R_{b}=$ 6.88 µm (and $>4$ µm) and the one between the second nearest-neighbors to be greater than $1.2R_{b}$. The optimal solutions are shown in the left column of Fig. 9.
@@ F.2 Local-detuning optimization to mitigate the tail interaction
For all geometrical configurations, we search for a pattern of local detuning $\epsilon_{v}$ to minimize $\delta\epsilon$. The optimal patterns are shown in the bottom inset of Fig. 20, and the expectation value of the domain-wall correlation operator $C(\bm{j})$ computed on the associated ground states is shown in the right column. In all cases, we achieve a fidelity $F\gtrsim 0.95$ between these ground states and the ones in the blockade limit, proving that the local adjustment of the detuning is an efficient strategy to mitigate the tail interaction.
@@ Appendix G Data analysis for the success probability
Figure 22: String occurrence of the raw data for the four configurations of the CWE setup (see Figs. 9 and 10). We report the occurrence of the most probable 10 strings divided by the number of measurements $N$. We mark the strings encoding the solution to the encoded MWIS problem with a star.
@@
Table 1: Post-processing for the success probabilities in Fig. 10. For the four configurations of the CWE setup, we report the number of measurements $N$ where the atoms are correctly loaded. We report the probability of measuring the strings encoding the solution to the MWIS problem when: considering all the measurements, denoted as $P_{\rm raw}$; discarding the strings that violate the blockade, denoted as $P_{\rm B}$.
@@
In the raw data, we observe both strings that violate the blockade constraint and strings with multiple domain walls per wire. We attribute these effects to a combination of factors, including non-adiabatic evolution, detection errors, and decoherence. To minimize these effects, we perform the following post-processing scheme. First, we discard the strings violating the blockade condition, obtaining a refined success probability, denoted as $P_{B}$ in Table 1. Then, we perform a greedy vertex addition strategy following Ref. [4], which yields the final refined success probability $P$, shown in the top-left of Fig. 10.
@@ Appendix H Ancillary atoms for the CWE gadget
Here, we discuss the extended two-domain-wall space obtained when modifying the CWE-gadget geometry via the introduction of ancillary atoms. The original two-domain-wall space [see Fig. 5(c)] is effectively extended at the non-logical boundary through the so-called copy states, as shown in Fig. 12. In this extended two-domain-wall space, we can control the diagonal terms on all the boundary states. This allows us to exactly satisfy the hopping-detuning balance condition in Eq. (9) on the entire two-domain-wall space, leading to the equal-superposition state as the ground state.
Particularly, we introduce the ancillary atoms $A_{w}$, $A_{Nw}$, and $B_{w}$ (with $w={x,y}$) to the CWE-gadget geometry, as shown in Fig. 12. We then parameterize the detuning and Rabi frequency on $A_{w}$ respectively as $\Delta+\epsilon_{A_{w}}$ and $\Omega_{A_{w}}$; on $A_{Nw}$ as $\Delta$ and $\Omega_{A_{Nw}}$; and on $B_{w}$ as $\epsilon_{B_{w}}$ and $\Omega_{B_{w}}$. Then, we work in the perturbative limit of $\epsilon_{i,w},\epsilon_{f,w},\epsilon_{A_{w}},\epsilon_{B_{w}}\ll\Delta$ and $\Omega,\Omega_{A_{w}},\Omega_{A_{Nw}},\Omega_{B_{w}}\ll\Delta$.
@@ H.1 Positioning of the ancillae $A$
The lowest energy subspace of $H_{0}$ is an extended two-domain-wall space. The states are now labeled by the coordinates of the domain walls and the occupation numbers $n_{A_{x}}$ and $n_{A_{y}}$ of $A_{x}$ and $A_{y}$ as $|j_{x},j_{y};n_{A_{x}},n_{A_{y}}\rangle$. For $n_{A_{x}}=n_{A_{y}}=0$, we retrieve the original two-domain-wall space; while, for $n_{A_{x}}=1$ and $n_{A_{y}}=0$ ($n_{A_{x}}=0$ and $n_{A_{y}}=1$), we retrieve the copy states of the non-logical boundary-line $\{j_{x}>\zeta_{x}+1,j_{y}=\zeta_{y}\}$ ($\{j_{x}=\zeta_{x},j_{y}>\zeta_{y}+1$}). The states located at the non-logical boundary-line $\{j_{x}>\zeta_{x}+1,j_{y}=\zeta_{y}\}$ ($\{j_{x}=\zeta_{x},j_{y}>\zeta_{y}+1\}$) have an additional copy state, characterized by the same values of $j_{x}$ and $j_{y}$ but with $n_{A_{x}}=1-n_{A_{y}}=1$ ($n_{A_{y}}=1-n_{A_{x}}=1$). At this point, the ancillae $A_{Nx}$ and $A_{Ny}$ do not further extend the lowest energy subspace, but they are needed in combination with the ancillae $B_{w}$, as explained in the next subsection. Indeed, the occupation numbers $n_{A_{Nx}}$ and $n_{A_{Ny}}$ are fixed to be the negation of $n_{A_{x}}$ and $n_{A_{y}}$, respectively: $A_{w}$ and $A_{Nw}$ are blockaded with each other, and the states with $n_{A_{w}}=n_{A_{Nw}}=0$ are energetically unfavorable.
In the lowest energy subspace of $H_{0}$, the dynamics are determined by the effective Hamiltonian in Eq. (13) with perturbation term given by ${H_{1,A}=H_{\epsilon}+H_{\epsilon,{\rm A}}+H_{\Omega}+H_{\Omega,\rm A}}$, where
@@
We can thus control the diagonal terms also along the non-logical boundary via the detuning on the ancillae. By setting ${\Omega_{A_{w}}=\Omega_{A_{Nw}}=\Omega}$, we homogeneously extend the hopping to the copy states and thus obtain the Hamiltonian represented in Fig. 12, with the addition of the orange sites only (and no blue sites).
Via this gadget modification, by setting $\epsilon_{A_{x}}=\epsilon_{A_{y}}=t$, we impose the hopping-detuning balance condition in Eq. (9) on the whole non-logical component of the two-domain-wall space boundary, apart from the internal corner states $|\zeta_{x}+2,\zeta_{y};1,0\rangle$ and $|\zeta_{x},\zeta_{y}+2;0,1\rangle$. This results in an incomplete delocalization of the ground-state wavefunction, motivating us to add further ancillary atoms as discussed below.

### v2

@@ E.1 Matrix-transfer method and bound state
Figure 17: Analysis of the bound state in a infinite $L$-shaped domain. (a) For $L\to\infty$, the domain of the Helmholtz equation in Eq. (6) acquires an infinite $L$ shape. The two leads are denoted by $1$ and $2$, with widths $\zeta_{1}$ and $\zeta_{2}$, and the inner region with “In”. Three different reference frames are used for the three regions. (b) Ground-state energy, $k^{2}_{\rm GS}$, computed via (black) finite-element method (FEM) and (orange) single-mode approximation (SMA) as a function of $r_{\zeta}=\zeta_{2}/\zeta_{1}$. The gap $d$ in Eq. (42) computed using FEM is shown in blue. Here, we set $L=40\zeta_{1}$.
Here, we generalize the matrix-transfer method presented in Ref. [84] to solve the Helmholtz equation with homogeneous DBC in a symmetric infinite $L$-shaped domain, to the asymmetric case (namely, leads of different widths). Following the notations of Fig. 17(a), we divide the Hilbert space into three distinct regions: two leads, denoted by $\Gamma\in\{1,2\}$, of different widths ($\zeta_{1}\equiv\zeta_{x}$ and $\zeta_{2}\equiv\zeta_{y}$); and an inner region, denoted by “In.” We employ three different reference frames: $\{x_{1},y_{1}\}$ in the lead extending along $y$, $\{x_{2},y_{2}\}$ in the lead extending along $x$, and $\{X,Y\}$ in the inner region.
@@
This equation implies ${M_{11}=0}$, from which we determine the value of $k$ for the bound state, $k_{\rm BS}$. We rescale the system by setting (without loss of generality) $\zeta_{1}=1$ and rescaling the energies accordingly. We find that $k_{\rm BS}$ satisfy the following equation:
@@ E.2 Energy gap scaling for arbitrary geometries
Here, we employ a finite-element method (FEM) to solve the Helmholtz equation with homogeneous DBC, in the finite $L$-shaped domain. For different values of $r_{\zeta}$, we compute how the energy gap ${\delta E=t\left(k^{2}_{\rm th}-k^{2}_{\rm GS}\right)}$ scales with $L$. We choose, without loss of generality, $L=1$ and numerically solve the equation rescaling the frame of coordinates as $\left\{x^{\prime}=x/L,y^{\prime}=y/L\right\}$ and the energies via ${k^{\prime}=k/L}$. We find the following scaling of the energy gap:
@@ E.3 Conformal mapping and bound state
The bound-state solution of the Helmholtz equation can be further investigated via a conformal mapping of the $L$-shaped domain into a unit disk. Here, the transformed equation admits a semianalytical treatment, which allows us to classify the eigenstates into bound states and propagating states at infinity. Without loss of generality, in the following discussion, we normalize the width of the two leads such that $\zeta_{1}\zeta_{2}=1$.
We parameterize the $L$-shaped domain $\mathcal{L}$ with a complex coordinate $w=\xi+i\eta$, and the unit disk ${\mathcal{D}=\{(x,y)|\sqrt{x^{2}+y^{2}}<1\}}$ with $z=x+iy$. We derive a conformal mapping between $w$ and $z$ in two steps, where we resort to an intermediate coordinate $u$ spanning the upper half of the complex plane. First, the mapping from $u$ to $w$ is achieved via a Schwarz-Christoffel transformation that sends the prevertices $u=-\zeta_{1}^{2},0,\zeta_{1}^{-2},\infty$ on the real axis $u$ to the vertices of the $L$-shaped polygon $w=+i\infty,0,+\infty,\zeta_{1}+i\zeta_{1}^{-1}$, respectively:
@@ E.3.1 Qualitative analysis of the eigenstates
The asymptotic behaviors of the eigenstates at infinity of $\mathcal{L}_{p}$ can be inferred from the analytical solution of Eq. (45) in the vicinity of the transformed points. The points at infinity in the two leads $\Gamma\in\{1,2\}$, namely, $w_{1}=-e^{-ip\pi}\infty$ and $w_{2}=+\infty$, are, respectively, mapped to $z_{1}=-e^{-i\varphi}$ and $z_{2}=e^{-i\varphi}$ on the unit disk. In the vicinity of these poles, Eq. (45) can be solved by separating the variables after switching to polar coordinates, i.e., $z-z_{\Gamma}=\rho e^{i\theta}$. To the lowest order in $\rho$, it results that $\chi_{p}\simeq\pi^{2}\zeta_{\Gamma}^{4p}/\rho^{2}$ has no angular dependence. By setting $\phi_{p}(\rho,\theta)=R(\rho)T(\theta)$, Eq. (45) returns ${T^{\prime\prime}+m^{2}T=0}$ and
@@
Here, the first inequality is derived from the spatial symmetries of $\phi_{p=0}$ and $\chi_{p}$. Indeed, the invariance of the operator $-\chi_{0}^{-1}\nabla^{2}$ upon the transformation $(x,y)\to(-x,-y)$ leads to $\phi_{p=0}^{2}(x,y)=\phi_{p=0}^{2}(-x,-y)$, and the Jacobians verify ${\chi_{p}(x,y)+\chi_{p}(-x,-y)>2\chi_{p=0}(x,y)}$ for any $p>0$ (using $\mu+1/\mu>2$ for $\mu>0,\mu\neq 1$). The second inequality stems from the variational principle for estimating the ground-state energy (i.e., $E_{p}[\phi]\geq E_{p}[\phi_{p}]$ for every $\phi$). Solving the Helmholtz equation directly in $\mathcal{L}_{0}$ results in $\phi_{p=0}=\frac{1-x^{2}-y^{2}}{\sqrt{[x^{2}+(y-1)^{2}][x^{2}+(y+1)^{2}]}}$ and $k_{p=0}^{2}=\pi^{2}$, and thus the ground-state energy in $\mathcal{L}_{p>0}$ is characterized by a strict upper bound: $k_{p>0}^{2}<\pi^{2}$.
@@
E.3.2 Semianalytical solution
We now present a possible semianalytical treatment of the ground state for $\mathcal{L}_{p}$ domains. In particular, the homotopy perturbation method (HPM) [96] can be applied to Eq. (45), leading to a recursive approximation of the bound state. Starting from an initial guess $\phi_{p,n=0}$, this method generates a sequence of un-normalized wavefunctions $\phi_{p,n}$ and corresponding energies $k^{2}_{p,n}$, via the following Poisson equation:
@@ Appendix F Analyses beyond the blockade limit
Here, we discuss how the analyses conducted in the blockade limit also hold when considering the Van der Waals interaction $U_{v,u}$ in Eq. (2) beyond the blockade radius. Specifically, we focus on the domain-wall localization phenomenon occurring in the CWE setup (see Secs. IVC and V), but we mention that analogous reasoning applies to the vertex wire and the crossing gadget.
Due to the tail interaction among nonblockaded atoms, the effective Hamiltonian in Eq. (5) acquires additional site-dependent diagonal terms. Specifically, the diagonal term $\epsilon(\bm{j})$ is given by the sum of the detuning $\Delta_{v}$ on the excited atoms of the domain-wall state $|\bm{j}\rangle$, plus the sum of the interactions $U_{v,u}$ between its pairs of excited atoms (at a distance $d_{v,u}$). In the case of the CWE setup of Fig. 1(d), due to its geometrical structure, these diagonal terms tend to localize the domain walls around the gadget. Indeed, to enforce the blockade constraints, the gadget atoms need to be placed at a shorter distance than those in the legs. This introduces an energetic penalty for states with more excited atoms near the gadget.
In the following, we focus on two possible strategies to mitigate this effect and recover the blockade-limit ground state with high fidelity. The first strategy relies on adjusting the position of the atoms along the vertex wires, while the second one relies on adjusting the local detuning on each atom. In both strategies, we aim to minimize the spreading of the diagonal terms $\delta\epsilon\equiv\max_{\bm{j}}{\left[\epsilon(\bm{j})\right]}-\min_{\bm{j}}{\left[\epsilon(\bm{j})\right]}$, since the blockade-limit is recovered in the $\delta\epsilon\ll t$ limit.
In the numerical simulations, we consider the parameters of Sec. VA: $\Omega=\Delta/3=7.5$ rad/µs, $C_{6}=5.42\times 10^{6}$ µm6 $\times$ rad/µs, $R_{b}=\left(C_{6}/\sqrt{(2\Omega)^{2}+\Delta^{2}}\right)^{1/6}=7.65$ µm, and the constraint of $d_{v,u}>4$ µm; include the tail interaction between nonblockade atoms; and employ the exact diagonalization method. In the implementation of Sec. VA, we adopted the position-based mitigation strategy.
@@ F.1 Position optimization to mitigate the tail interaction
Here, we optimize the position of the vertex wire atoms to minimize $\delta\epsilon$, while fixing the position of the five atoms at the intersection of the two vertex wires. In particular, we set the distance between $G_{c}$ and its neighbors [see inset in Fig. 1(d)] to $0.9R_{b}/\sqrt{2}$ and the other distances (e.g., between $G_{x}$ and $G_{y}$) to $0.9R_{b}$. Then, we search for an optimal position pattern along the vertex wires, constraining the distance between neighboring atoms to be smaller than $6.880.9R_{b}=$ 6.88 µm (and $>4$ µm) and the one between the second nearest neighbors to be greater than $1.2R_{b}$. The optimal solutions are shown in the left column of Fig. 9.
@@ F.2 Local-detuning optimization to mitigate the tail interaction
[…]
@@ Appendix G Data analysis for the success probability
Figure 22: String occurrence of the raw data for the four configurations of the CWE setup (see Figs. 9 and 10). We report the occurrence of the most probable ten strings divided by the number of measurements $N$. We mark the strings encoding the solution to the encoded MWIS problem with a star.
@@
Table 2: Postprocessing for the success probabilities in Fig. 10. For the four configurations of the CWE setup, we report the number of measurements $N$ where the atoms are correctly loaded. We report the probability of measuring the strings encoding the solution to the MWIS problem when: considering all the measurements, denoted as $P_{\rm raw}$; discarding the strings that violate the blockade, denoted as $P_{\rm B}$.
@@
In the raw data, we observe both strings that violate the blockade constraint and strings with multiple domain walls per wire. We attribute these effects to a combination of factors, including nonadiabatic evolution, detection errors, and decoherence. To minimize these effects, we perform the following postprocessing scheme. First, we discard the strings violating the blockade condition, obtaining a refined success probability, denoted as $P_{B}$ in Table 2. Then, we perform a greedy vertex addition strategy following Ref. [4], which yields the final refined success probability $P$, shown in the top left of Fig. 10.
@@ Appendix H Ancillary atoms for the CWE gadget
Here, we discuss the extended two-domain-wall space obtained when modifying the CWE-gadget geometry via the introduction of ancillary atoms. The original two-domain-wall space [see Fig. 5(c)] is effectively extended at the nonlogical boundary through the so-called copy states, as shown in Fig. 12. In this extended two-domain-wall space, we can control the diagonal terms on all the boundary states. This allows us to exactly satisfy the hopping-detuning balance condition in Eq. (9) on the entire two-domain-wall space, leading to the equal-superposition state as the ground state.
Particularly, we introduce the ancillary atoms $A_{w}$, $A_{Nw}$, and $B_{w}$ (with $w={x,y}$) to the CWE-gadget geometry, as shown in Fig. 12. We then parameterize the detuning and Rabi frequency on $A_{w}$, respectively, as $\Delta+\epsilon_{A_{w}}$ and $\Omega_{A_{w}}$; on $A_{Nw}$ as $\Delta$ and $\Omega_{A_{Nw}}$; and on $B_{w}$ as $\epsilon_{B_{w}}$ and $\Omega_{B_{w}}$. Then, we work in the perturbative limit of $\epsilon_{i,w},\epsilon_{f,w},\epsilon_{A_{w}},\epsilon_{B_{w}}\ll\Delta$ and $\Omega,\Omega_{A_{w}},\Omega_{A_{Nw}},\Omega_{B_{w}}\ll\Delta$.
@@ H.1 Positioning of the ancillae $A$
The lowest-energy subspace of $H_{0}$ is an extended two-domain-wall space. The states are now labeled by the coordinates of the domain walls and the occupation numbers $n_{A_{x}}$ and $n_{A_{y}}$ of $A_{x}$ and $A_{y}$ as $|j_{x},j_{y};n_{A_{x}},n_{A_{y}}\rangle$. For $n_{A_{x}}=n_{A_{y}}=0$, we retrieve the original two-domain-wall space; while, for $n_{A_{x}}=1$ and $n_{A_{y}}=0$ ($n_{A_{x}}=0$ and $n_{A_{y}}=1$), we retrieve the copy states of the nonlogical boundary line $\{j_{x}>\zeta_{x}+1,j_{y}=\zeta_{y}\}$ ($\{j_{x}=\zeta_{x},j_{y}>\zeta_{y}+1$}). The states located at the nonlogical boundary line $\{j_{x}>\zeta_{x}+1,j_{y}=\zeta_{y}\}$ ($\{j_{x}=\zeta_{x},j_{y}>\zeta_{y}+1\}$) have an additional copy state, characterized by the same values of $j_{x}$ and $j_{y}$ but with $n_{A_{x}}=1-n_{A_{y}}=1$ ($n_{A_{y}}=1-n_{A_{x}}=1$). At this point, the ancillae $A_{Nx}$ and $A_{Ny}$ do not further extend the lowest-energy subspace, but they are needed in combination with the ancillae $B_{w}$, as explained in the next subsection. Indeed, the occupation numbers $n_{A_{Nx}}$ and $n_{A_{Ny}}$ are fixed to be the negation of $n_{A_{x}}$ and $n_{A_{y}}$, respectively: $A_{w}$ and $A_{Nw}$ are blockaded with each other, and the states with $n_{A_{w}}=n_{A_{Nw}}=0$ are energetically unfavorable.
In the lowest-energy subspace of $H_{0}$, the dynamics are determined by the effective Hamiltonian in Eq. (13) with perturbation term given by ${H_{1,A}=H_{\epsilon}+H_{\epsilon,{\rm A}}+H_{\Omega}+H_{\Omega,\rm A}}$, where
@@
We can thus control the diagonal terms also along the nonlogical boundary via the detuning on the ancillae. By setting ${\Omega_{A_{w}}=\Omega_{A_{Nw}}=\Omega}$, we homogeneously extend the hopping to the copy states and thus obtain the Hamiltonian represented in Fig. 12, with the addition of the orange sites only (and no blue sites).
Via this gadget modification, by setting $\epsilon_{A_{x}}=\epsilon_{A_{y}}=t$, we impose the hopping-detuning balance condition in Eq. (9) on the whole nonlogical component of the two-domain-wall space boundary, apart from the internal corner states $|\zeta_{x}+2,\zeta_{y};1,0\rangle$ and $|\zeta_{x},\zeta_{y}+2;0,1\rangle$. This results in an incomplete delocalization of the ground-state wavefunction, motivating us to add further ancillary atoms as discussed below.